# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v1

@@ Appendix F Results on the E-MNIST dataset
In order to analyze the scalability of ProbSENN in tasks consisting of larger number of classes, we evaluated its performance in the E-MNIST dataset Cohen et al. [2017] (47 classes), an extension of MNIST which includes both numeric and alphabetic characters. As it can be seen in Figure 17, ProbSENN is still capable of learning suitable distribution over the prototypes for the 47 classes, yielding realistic and diverse prototypes, while achieving a test accuracy of 0.85.
@@
Figure 17: Prototypes sampled for the E-MNIST dataset (47 classes), consisting of handwritten representations of numbers and letters.
Appendix G Harnessing uncertainty to improve predictive accuracy and reliability: Results
Table 2: Predictive accuracy percentages (mean and standard deviation for 10 runs) obtained by Prob-PSENN for the following configurations: using the MLP-Autoencoder, using the CNN-Autoencoder, including elastic deformation as data augmentation (AUG), and discarding the inputs with an uncertainty rate $\mathscr{U}_{E}(x)$ (see Equation 4 in the paper) higher than a threshold $\alpha$ (UNC).
Configuration MNIST F-MNIST K-MNIST
Prob-PSENN ($l=10$, MLP) 97.59 $\pm$ 0.48 88.77 $\pm$ 0.50 88.92 $\pm$ 1.00
 + UNC ($\alpha=0.99$) 98.99 $\pm$ 0.42 89.16 $\pm$ 0.52 96.19 $\pm$ 1.42
 + UNC ($\alpha=0.95$) 99.68 $\pm$ 0.13 90.52 $\pm$ 0.67 98.77 $\pm$ 0.39
Prob-PSENN ($l=10$, CNN) 98.28 $\pm$ 0.17 89.47 $\pm$ 0.34 89.06 $\pm$ 0.34
 + UNC ($\alpha=0.99$) 98.83 $\pm$ 0.25 89.52 $\pm$ 0.41 89.58 $\pm$ 0.50
 + UNC ($\alpha=0.95$) 99.68 $\pm$ 0.09 90.22 $\pm$ 0.59 94.58 $\pm$ 0.74
Prob-PSENN ($l=10$, MLP + AUG) 98.53 $\pm$ 0.08 87.78 $\pm$ 0.50 93.17 $\pm$ 0.33
 + UNC ($\alpha=0.99$) 98.97 $\pm$ 0.14 87.81 $\pm$ 0.49 95.29 $\pm$ 0.65
 + UNC ($\alpha=0.95$) 99.67 $\pm$ 0.05 88.29 $\pm$ 0.57 98.44 $\pm$ 0.33
Prob-PSENN ($l=10$, CNN + AUG) 98.75 $\pm$ 0.17 87.98 $\pm$ 0.46 89.40 $\pm$ 1.14
 + UNC ($\alpha=0.99$) 99.00 $\pm$ 0.18 88.00 $\pm$ 0.46 89.56 $\pm$ 1.09
 + UNC ($\alpha=0.95$) 99.74 $\pm$ 0.09 88.45 $\pm$ 0.52 93.53 $\pm$ 1.28
Table 3: Percentage of discarded inputs when filtering by uncertainty (mean and standard deviation for 10 runs).
Threshold Configuration MNIST F-MNIST K-MNIST
$\alpha=0.99$ Prob-PSENN ($l=10$, MLP) 2.77 $\pm$ 0.53 0.79 $\pm$ 0.26 11.49 $\pm$ 2.31
Prob-PSENN ($l=10$, CNN) 1.01 $\pm$ 0.26 0.13 $\pm$ 0.10   0.70 $\pm$ 0.40
Prob-PSENN ($l=10$, MLP + AUG) 0.90 $\pm$ 0.14 0.05 $\pm$ 0.03   4.25 $\pm$ 0.97
Prob-PSENN ($l=10$, CNN + AUG) 0.52 $\pm$ 0.11 0.04 $\pm$ 0.04   0.17 $\pm$ 0.11
$\alpha=0.95$ Prob-PSENN ($l=10$, MLP) 7.69 $\pm$ 0.50 5.16 $\pm$ 0.86 24.89 $\pm$ 1.33
Prob-PSENN ($l=10$, CNN) 5.27 $\pm$ 0.21 2.57 $\pm$ 0.45 12.38 $\pm$ 1.10
Prob-PSENN ($l=10$, MLP + AUG) 5.58 $\pm$ 0.23 1.84 $\pm$ 0.29 17.02 $\pm$ 1.01
Prob-PSENN ($l=10$, CNN + AUG) 4.67 $\pm$ 0.21 1.75 $\pm$ 0.30    9.59 $\pm$ 1.17
Table 4: Error rate (%) of the deterministic baseline model on the inputs receiving high/low uncertainty by Prob-PSENNs (with $l=10$ and an MLP-Autoencoders).
Threshold Configuration MNIST F-MNIST K-MNIST
$\alpha=0.99$ High uncert. ($\mathscr{U}_{E}(x)>\alpha$) 13.10 $\pm$ 1.87 25.56 $\pm$ 6.72 37.12 $\pm$ 2.69
Low uncert.  ($\mathscr{U}_{E}(x)\leq\alpha$)   0.73 $\pm$ 0.13   9.73 $\pm$ 0.32   5.36 $\pm$ 0.51
$\alpha=0.95$ High uncert. ($\mathscr{U}_{E}(x)>\alpha$)   8.75 $\pm$ 1.24 25.55 $\pm$ 3.29 27.12 $\pm$ 1.38
Low uncert.  ($\mathscr{U}_{E}(x)\leq\alpha$)   0.41 $\pm$ 0.09   8.99 $\pm$ 0.35   2.92 $\pm$ 0.33
@@ Appendix H Broader Impact
As specified above, the goal of Prob-PSENNs is to promote a transparent-by-design DNN architecture while maintaining competitive predictive performance. In addition to transparency, Prob-PSENNs provide tools to quantify the uncertainty in the model’s predictions and explanations, which further enhances its safety and reliability. Therefore, our work is aligned with three important needs of deep learning models: interpretability, transparency, and reliability. However, a current concern in deep learning models is their lack of robustness to adversarial attacks Yuan et al. [2019], Szegedy et al. [2014], Biggio et al. [2013], a vulnerability that also threatens explanation methods and self-explainable models Vadillo et al. [2021], Aivodji et al. [2019], Lakkaraju and Bastani [2020], Ghorbani et al. [2019]. Therefore, in order to ensure a safe and responsible deployment of Prob-PSENNs in practice, it remains crucial to carefully assess the implications of adversarial attacks, and, consequently, to evaluate the robustness of the model against them.

### v2

@@ Appendix G Results on the E-MNIST dataset
In order to analyze the scalability of Prob-PSENN in tasks consisting of larger number of classes, we evaluated its performance in the E-MNIST dataset [9] (47 classes), an extension of MNIST which includes both numeric and alphabetic characters. As it can be seen in Figure G.22, Prob-PSENN is still capable of learning suitable distribution over the prototypes for the 47 classes, yielding realistic and diverse prototypes, while achieving a test accuracy of 0.89.
@@
Figure G.22: Prototypes sampled for a Prob-PSENN trained on the E-MNIST [9] dataset (47 classes), consisting of handwritten representations of numbers and letters.
@@ Appendix H Broader Impact
As specified above, the goal of Prob-PSENNs is to promote a transparent-by-design DNN architecture while maintaining competitive predictive performance. In addition to transparency, Prob-PSENNs provide tools to quantify the uncertainty in the model’s predictions and explanations, which further enhances its safety and reliability. Therefore, our work is aligned with three important needs of deep learning models: interpretability, transparency, and reliability. However, a current concern in deep learning models is their lack of robustness to adversarial attacks [45, 41, 6], a vulnerability that also threatens explanation methods and self-explainable models [42, 2, 25, 14]. Therefore, in order to ensure a safe and responsible deployment of Prob-PSENNs in practice, it remains crucial to carefully assess the implications of adversarial attacks, and, consequently, to evaluate the robustness of the model against them.